OpenStudy (anonymous):

A solid has a circular base of radius 3. If every plane cross section perpendicular to the x- axis is an equilateral triangle, the. It's volume is..

OpenStudy (experimentx):

??? couldn't picture at all

OpenStudy (anonymous):

*then, it's volume is ...

OpenStudy (turingtest):

imagine a bunch of triangles popping out in the z-direction

OpenStudy (dumbcow):

circle: x^2 + y^2 = 9 --> y = sqrt(9-x^2)  volume is sum of all the areas of the triangles area = sqrt3/4 *(2y)^2 = sqrt3 y^2 = sqrt3(9-x^2) \[V = \sqrt{3}\int\limits_{-3}^{3}(9-x^{2}) dx\]